OpenStudy (anonymous):

Please help me with this identity: tan ^2(x)/(1+tan^2(x))=sin^2(x)

OpenStudy (anonymous):

\[\frac{ \tan ^{2}x }{ 1+\tan ^{2}x } = \sin ^{2}x\]

OpenStudy (anonymous):

break it out, the denominator = sec^2(x) = 1/ cos^2 x

OpenStudy (anonymous):

the numerator is sin^2/cos^2 . combine them, you have the result. hope this helps

OpenStudy (anonymous):

thanks, i've been doing this all morning and didn't even realize there was a pythagorian identity right there

OpenStudy (anonymous):

hihi,...glad you got